Simplify.Express in simplest radical form
(√3-√5)(√3-√5)

Answer by algebrahouse.com(1659)

(√3 - √5)(√3 - √5)
= √3(√3) + √3(-√5) - √5(√3) - √5(√5) {used foil method}
= √9 - √15 - √15 - √25 {multiplied through}
= 3 - 2√15 - 5 {evaluated square roots and combined like terms}
= -2 - 2√15 {combined like terms}

For more help from me, visit:
www.algebrahouse.com